A rectangular has a perimeter of 62 centimeters. If the sum of the length and twice the width is 42 centimeters, find the length.

Answer by Cromlix(4381)

Length = x
Width = y
Perimeter = 2x + 2y

2x + 2y = 62
x + 2y = 42
Subtract
x = 20
Substitute x = 20
into x + 2y = 42
20 + 2y = 42
2y = 42 - 20
2y = 22
y = 11
Length = 20cm
Width = 11cm
Hope this helps.
:-)